### v1

# Surrogate normal-forms for the numerical bifurcation and stability analysis of Navier-Stokes flows via machine learning

Abstract
Inspired by the Equation-Free multiscale modelling approach, we show how the “embed-learn-lift” framework can be exploited for the construction of surrogate “normal-forms” (reduced-order models of minimal dimensionality) of high-fidelity Navier-Stokes simulations, that can in turn be used for computationally efficient and numerically accurate bifurcation and stability analysis tasks. The framework consists of four main steps. In the first step, we apply manifold learning to discover the intrinsic dimension of the emergent latent space of the complex, high-dimensional Navier-Stokes spatio-temporal dynamics over the parameter space. In the second stage, we construct “normal-forms” like reduced-order models (ROMs)-in the sense that we capture the correct dimension for describing the emergent dynamics- on the latent space using Gaussian Process regression (GPR). Then, based on the constructed surrogate “normal-form” models, which dramatically reduce computational complexity, we exploit the toolkit of numerical bifurcation analysis theory to construct the bifurcation diagrams in the latent spaces, and perform systematic stability analysis. This allows, for example, the continuation of branches of limit cycles emanating from Andronov-Hopf bifurcations; such an analysis is intractable for high-dimensional systems such as the Navier-Stokes equations using standard Jacobian-full based computations. Finally, by solving the pre-image problem, we reconstruct the bifurcation diagrams in the original high-dimensional space. The proposed methodology is demonstrated by focusing on two suitable test-bed flow configurations: the wake flow past an infinitely long circular cylinder and the planar sudden-expansion channel flow. These two flow systems exhibit Andronov-Hopf and pitchfork (symmetry-breaking) bifurcations by increasing the Reynolds number $Re$, respectively. The proposed methodology successfully identifies the appropriate dimensionality of the latent space of the Navier-Stokes equations—and based on them constructs ROMs referred to here as "surrogate normal-forms" using Gaussian Process regression (GPR)—that enable the accurate tracing of the bifurcating solutions and their stability analysis. This includes the accurate tracing of limit cycles, their period, and their stability characteristics, as determined through the computation of Floquet multipliers.
1 Introduction
Numerical simulations of nonlinear phenomena in fluid mechanics have reached today an unprecedented level of detail thanks to the continued advancement of computational algorithms and hardware resources. The complex multiscale spatio-temporal behavior exhibited by fluid flows can be readily reproduced using high-fidelity models, allowing for accurate high-dimensional descriptions of the physical mechanisms at play [1].
However, understanding the mechanisms that govern complex phenomena such as flow instabilities and turbulence, requires more than time-marching simulations: while modern Computational Fluid Dynamics (CFD) codes efficiently integrate transients and even find steady states via Newton solvers, they cannot reliably locate precisely the critical (tipping) points at which instabilities appear or trace the resulting unstable branches. For a systematic and precise analysis of dynamical behaviors, particularly the emergence and evolution of oscillatory patterns (even unstable ones that can explain the mechanisms that lead to turbulence), a systematic bifurcation analysis is indispensable. The numerical bifurcation theory toolkit offers a suite of algorithms and software tools—such as AUTO [2] and MATCONT [3, 4]—that facilitate the continuation of both stable and unstable steady states, as well as limit cycles and critical bifurcation points. These packages have been built on decades of dedicated research and development, incorporating extensive expertise in nonlinear dynamics and numerical continuation methods. Their robust algorithms and longstanding validation make them trusted choices for performing accurate bifurcation analysis and continuation, including Hopf, fold, and period-doubling bifurcations, as well as codimension-2 bifurcation tracking for low- to medium-scale systems. However, for high-fidelity CFD simulations, storing and factorizing high-dimensional matrices renders tasks such as limit-cycle continuation computationally intractable beyond small-scale test cases [5]. Matrix-free Krylov subspace methods, such as Newton–Krylov solvers, offer a scalable approach for large scale partial differential equation (PDE) systems such as those resulting in fluid flows by avoiding the explicit formation of Jacobians and leveraging efficient matrix-vector products [6, 5, 7, 8, 9, 10]. In particular, for the Navier-Stokes equations, Newton–Krylov continuation of periodic orbits has been performed in [6]. However, these methods lack the comprehensive algorithmic capabilities of the aforementioned established bifurcation packages. Such tools provide robust and state-of-the-art techniques for detecting and continuing bifurcations, and offer user-friendly interfaces for systematic analysis. In contrast, matrix-free approaches often require custom implementation and may not support advanced analysis as codimension-2 bifurcation tracking, thus limiting their applicability for in-depth bifurcation studies.
Reduced-order models (ROMs) in CFD  [11, 12, 13, 14, 15, 16, 17, 1] reduce the computational cost and memory demands by orders of magnitude, allowing for faster simulations that can capture qualitatively the emergent behaviour, and the systematic bifurcation analysis that with the high-fidelity full-order models would be overwhelming difficult to perform. The fundamental idea behind ROMs is that, despite a complex nature due to the inherent nonlinear behavior, fluid flows often exhibit a few dominant coherent structures, which contain coarse but valuable information about the underlying dynamics over the whole parameter space. A prominent example is the derivation of the Lorenz ordinary differential equations (ODEs) resulting from a low-dimensional truncation of the Navier–Stokes–Boussinesq PDE governing Rayleigh–Bénard convection. To derive such models, one can explicitly enforce knowledge of the high-dimensional governing equations, thus obtaining the so-called intrusive ROMs. A prominent paradigm of ROMs is represented by the so-called normal-forms, namely the simplest possible ROMs of complex/high-dimensional dynamical systems around a bifurcation point [18]. They are constructed by systematically truncating nonessential higher-order dimensions and terms, thus allowing for a better understanding of the intrinsic dimension of the emergent dynamics, greatly reducing computational cost and facilitating various numerical analysis and control tasks. Normal-forms are of vast importance in CFD because they allow complex, high-dimensional fluid equations—such as the Navier–Stokes system discretized over vast grids—to be rigorously reduced near critical parameter values (e.g., a critical Reynolds number) onto a low-dimensional model, greatly reducing computational cost. This reduction enables a deeper insight into how and why bifurcations emanate in complex fluid systems, rendering numerical continuation, stability analyses, and the design of controllers easier.
The traditional way for deriving such ROMs for fluid flows starts by linearizing the full Navier–Stokes equations around a base state at the bifurcation point and identifying the few critical eigenmodes whose growth rates are near zero. One then uses the invariant/center manifold theory [19, 11, 20, 21, 22, 23, 24] or spectral sub-manifold theory [25, 26, 27] to parameterize a locally invariant, low-dimensional manifold tangent to those modes—expressed usually as a power-series, thus truncating higher-order terms to obtain a self-consistent reduced ODE system that captures the key nonlinear dynamics around the bifurcation point. The theory of approximate inertial manifolds [12, 11] has also been used for the Navier-Stokes equation for the construction of a finite-dimensional invariant manifold onto which all trajectories are exponentially attracted, effectively capturing the global long-term dynamics of the flow. This reduces the infinite-dimensional PDE to a finite set of ODEs that can approximate the global dynamics.
On the other side, data-driven methods have been propelled in recent years by the exponential growth of machine learning (ML) and deep learning methodologies, which aim to construct surrogate ROMs from high-fidelity spatio-temporal simulations [28, 29, 30, 16]. The first step involves projecting the high-dimensional spatio-temporal flow dynamics onto a carefully chosen (low-dimensional) latent space. This in CFD is typically done via manifold learning techniques like Proper Orthogonal Decomposition (POD) [31, 32, 33, 34]. POD offers a direct, closed-form solution to the pre-image problem, i.e., the reconstruction of the CFD dynamics back in the original high-dimensional space. Nonlinear manifold learning algorithms, such as diffusion maps (DMs), have also been exploited  [35, 36, 37, 38, 39, 40, 41, 42] to deal with more complex dynamics, including chaotic flows [42]. In these approaches, in contrast to the straightforward closed-form mapping between full and reduced spaces provided by POD, the solution of the pre-image problem suffers from ill-posedness, and different techniques such as geometric harmonics and kernel ridge regression have been proposed (for a review, see [43, 44]). Machine learning-based architectures, such as autoencoders, have also been used for learning a set of coordinates that can parametrize the latent space dynamics [45, 46, 47, 48, 49]. On the part of the construction of surrogate ROMs via machine learning, various approaches have been proposed including Sparse identification of nonlinear dynamical systems (SINDy) [50, 51, 52, 53] coupled also with auto-encoders  [54, 55], feedforward/recurrent neural networks [56, 57, 58, 39, 59, 40, 60, 61, 62], cluster-based networks [63], spectral submanifolds [64] and gaussian process regression models (GPR) [65, 66, 67, 68]. For example, SINDy has been used to model a range of fluid flows, including laminar and turbulent wakes [51], convective [52] and shear [53] flows, and it has been also applied in combination with auto-encoders to learn latent space coordinates before finding the dynamical system to evolve [54]. Deep Neural Network-based Auto Encoders (DNN-AE) have been recently employed to identify reduced-order coordinates of a turbulent flow over a building-like structure on a nonlinear manifold [48]. Diffusion maps and neural networks have been leveraged to build predictive reduced-order models of a vertically falling liquid film flow in the case of sparse data [40]. A combination of $\beta$-variational autoencoders and transformers has been used to learn parsimonious and near-orthogonal reduced-order models for two-dimensional viscous flows in both periodic and chaotic regimes [69]. In [70, 71, 72, 73, 74], ROMs and bifurcation diagrams of equilibria are built from steady-state snapshots of Navier–Stokes simulations, targeting “stable bifurcating branches” using intrusive POD Galerkin projections with applications to triangular lid-driven cavity and sudden-expansion channel flows. This approach has also been used for the control of solutions on the steady-state unstable branch [73] and generalized in cases of fluid-structure interaction problems [74].
Here, building on previous efforts rooted in the celebrated Equation-Free multiscale framework [75], we propose a four-stage full data-driven framework for the construction of surrogate “normal-forms” via machine learning that can approximate not only qualitatively (as the traditional normal-form models) but also quantitatively, with high accuracy, the emergent dynamics of the Navier-Stokes equations. At this point, we note that we use the term “normal-forms” to emphasize the (main) challenge in complex systems modelling, that of learning the correct dimensionality of the latent dynamics across the parameter space, rather than reproducing the specific ROMs with minimal polynomial structures or local truncations as meant in local bifurcation theory. Consequently, our “normal-forms” for the Navier-Stokes equations represent ROMs with the correct dimensionality needed to approximate quantitatively the dynamics beyond the limited vicinity of the bifurcation points, which classical normal-forms cannot. Based on the fully data-driven constructed surrogate models, we then exploit the power of state-of-the-art continuation toolboxes such as MATCONT [3, 76] and AUTO [2], to perform computationally efficient and highly accurate bifurcation and stability analysis of the Navier-Stokes equations. Hence, we dramatically reduce computational and memory demands, yet still allow all the state-of-the-art bifurcation and stability (continuation of solution branches, even of unstable steady-states, continuation of limit cycles and their exact period, branch switching, the calculation of eigenvalues and Floquet multipliers, continuation of codimension-two bifurcations, continuation of homoclinic and heteroclinic bifurcations, Bautin and Takens–Bogdanov [77, 78, 79, 80, 81, 82, 83]). Such tasks would be intractable in the high-dimensional space generated by the full Navier–Stokes simulator, where one has to rely on linear (or weakly nonlinear) stability analysis of equilibria and on local theoretical normal-forms to trace bifurcation diagrams [83].
For our illustration, the proposed methodology is demonstrated by constructing “normal-forms” for two suitable test-bed flow configurations: the wake flow past an infinitely long circular cylinder and the planar sudden-expansion channel flow. These two flow systems exhibit Andronov-Hopf and pitchfork (symmetry-breaking) bifurcations by increasing the Reynolds number $Re$, respectively. In the first step, we apply manifold learning to learn the intrinsic dimensionality of complex high-dimensional flow fields over the parameter space generated by Navier–Stokes direct numerical simulations. In the second stage, we construct surrogate global normal-forms in the latent space using Gaussian Process regression (GPR) models. Then, based on the constructed global normal-forms, we exploit the toolkit of numerical bifurcation analysis theory (here using MATCONT) to construct the bifurcation diagrams in the low-dimensional latent spaces. Finally, by solving the pre-image problem, we reconstruct the bifurcation diagrams in the original high-dimensional space. The proposed methodology enables the identification of a single surrogate accurate “normal-form” that spans the entire range of the bifurcation parameter, allowing for the accurate reconstruction of the full bifurcation diagram, including the continuation of limit cycles and the corresponding stability analysis.
The rest of the work is organized as follows. In Section 2, we describe the manifold learning-GPR framework developed to learn data-driven ROMs suitable for the bifurcation analysis task. The two flow configurations employed to demonstrate the methodology are illustrated in Section 3, and results are discussed in Section 4. Conclusions are finally provided in Section 5.
2 Methodology
[…]
Figure 1: Four-stage data-driven framework for bifurcation and stability analysis of fluid flows in latent spaces: (a) projection of high-dimensional Navier-Stokes solutions to the latent space (manifold learning); (b) modelling of the latent dynamics via Gaussian Process regression; (c) computation of the latent bifurcation diagram and stability analysis; (d) reconstruction of the physical space solutions (pre-image problem).
The pipeline—rooted to the celebrated Equation-Free (EF) mutliscale numerical analysis framework [75]—for the systematic data-driven bifurcation and stability analysis of the Navier-Stokes equations consists of four main steps (see Fig. 2): (a) use manifold learning to “restrict/embed” the high-dimensional spatio-tempotal patterns into appropriate latent spaces, (b) learning surrogate ROMs in these latent spaces (here using Gaussian Regression Processes, i.e. GPR), (c) performing bifurcation and stability analysis in the latent spaces, (d) “lifting” to the high-dimensional space. A key distinction between our proposed framework and the traditional EF approach that was also coupled with POD for incompressible flows [84] is that, in the latter, one constructs local mappings and computes the quantities needed for bifurcation analysis via short bursts of microscopic simulations to bridge detailed simulations and emergent dynamics (captured by a few moment variables). In contrast, the proposed method is an extension to the traditional EF framework, in the sense that it learns a global model via machine learning from long-term, high-fidelity simulations spanning the entire parameter space. This global perspective enables the use of advanced nonlinear manifold-learning techniques, such as diffusion maps, and techniques for the solution of the pre-image problem, such as geometric harmonics or autoencoders, to construct maps between the high-dimensional and the latent space for complex dynamics like turbulence. As we have demonstrated in a series of works [57, 39, 85, 41, 61, 86], this framework can accurately reproduce the emergent bifurcation diagrams for simple one-dimensional PDEs and time series. Here, the framework is exploited to deal with the rich spatio-temporal complexity of the Navier–Stokes equations, representing a significant advance. We show for the first time that the proposed pipeline efficiently uncovers a low-dimensional representation that can accurately approximate the emergent dynamics via global surrogate models, whose governing equations closely resemble normal-forms of the correct dimension. This in turn allows the use of the state-of-the-art numerical bifurcation analysis toolkit, ensuring fidelity across a wide range of Reynolds numbers and the execution of tasks such as the continuation of limit cycles and their period, as well as their stability analysis. The latter task is done for the first time and showcases the potential of the framework to perform advanced bifurcation analysis. For the broader CFD community, the framework signals a promising new direction: the construction of fully data-driven normal-form ROMs, offering both high accuracy and efficiency, reducing computational expense by orders of magnitude. This will facilitate the execution of important tasks in CFD going beyond bifurcation analysis, such as real-time optimization and control.
Here, for our illustrations, we consider the incompressible 2D Navier–Stokes partial differential equations in the dimensionless form:
$\dfrac{\partial u}{\partial x}+\dfrac{\partial v}{\partial y}$ $=$ $0,$ (1a)
$\dfrac{\partial u}{\partial t}+u\dfrac{\partial u}{\partial x}+v\dfrac{\partial u}{\partial y}$ $=$ $-\dfrac{1}{\rho}\dfrac{\partial p}{\partial x}+\dfrac{1}{Re}\bigg{(}\dfrac{\partial^{2}u}{\partial x^{2}}+\dfrac{\partial^{2}u}{\partial y^{2}}\bigg{)},$ (1b)
$\dfrac{\partial v}{\partial t}+u\dfrac{\partial v}{\partial x}+v\dfrac{\partial v}{\partial y}$ $=$ $-\dfrac{1}{\rho}\dfrac{\partial p}{\partial y}+\dfrac{1}{Re}\bigg{(}\dfrac{\partial^{2}v}{\partial x^{2}}+\dfrac{\partial^{2}v}{\partial y^{2}}\bigg{)}.$ (1c)
where the variables $u\equiv u(x,y,t)$ and $v\equiv v(x,y,t)$ represent the streamwise and normal-to-flow velocity components respectively, across the 2D spatial domain ($x$, $y$) over time $t$. The variable $p\equiv p(x,y,t)$ denotes the pressure, and $\rho$ stands for the (constant) density. The governing bifurcation parameter is the Reynolds number $Re$, which represents the relative importance between inertia and viscous effects within the flow. The dimensional variables involved in the definition of $Re$ are case-specific; for the flow configurations considered in this work, $Re$ is given in Eq. (21) and Eq. (23) of the following Section 3.
The representational fidelity of ROMs for accurately describing high-dimensional systems, like the Navier-Stokes equations (1a,1b, 1c), relies on the key assumption that the long-term dynamics evolve on a low-dimensional inertial manifold. This manifold is a finite-dimensional, attracting, and invariant subset [87, 88, 89] embedded in the full-order model’s infinite-dimensional phase space. In practice, ROMs are constructed by identifying approximate inertial manifolds (AIMs), which can be parameterized using a small number of dominant modes. A paradigmatic example is the celebrated Lorenz equations, derived as a three-mode reduction of the Rayleigh-Bénard convection equations, which retains the chaotic attractor characteristics of the full-order PDE. Data-driven AIM parameterizations can be constructed numerically using, for example, POD embeddings or nonlinear manifold learning embeddings, as recently shown [42], but also autoencoders [49], for highly nonlinear phenomena. Such data-driven algorithms complement analytical AIM theory by providing latent coordinates for the slow/long-term dynamics, enabling reduced modeling of complex dissipative systems from data. For a detailed discussion between data-driven ROMs and AIMs, see [49].
2.1 The four-stage proposed data-driven framework
Our aim, to first construct a non-intrusive dynamical ROM of minimal dimensions, reflecting a normal-form-ROM, by assuming a given set $X\in\mathbb{R}^{N\times m}$ of $M$ spatio-temporal snapshots $\{\mathbf{u}_{m}\}_{m=1,\ldots,M}$ that arise from the discretized Navier-Stokes equations, or from fluid flow experiments. Each snapshot at time $t_{m}$ contains the velocity variables across the spatial domain; i.e., $\mathbf{u}_{m}=[u(x_{i},y_{j},t_{m}),v(x_{i},y_{j},t_{m})]^{\top}\in\mathbb{R}^{2N_{g}}$, where $i=1,\ldots,N_{x}$, $j=1,\ldots,N_{y}$ and $N_{g}=N_{x}\times N_{y}$ is the number of grid points. Based on the above rationale, we furthermore assume that these points lie on a smooth $d$-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, and we seek a data-driven mapping:
$\varPhi:X\;\longrightarrow\;A\subset\mathbb{R}^{d},\quad d\ll N,$ (2)
such that $\mathbf{a}_{m}=\varPhi(\mathbf{u}_{m})\in\mathbb{R}^{d}$ captures the dominant coherent structures and emergent dynamics of the flow, where $d$ is an approximation of the intrinsic dimension of $\mathcal{M}$. The construction of this mapping is the first stage of the proposed framework, stage (a) in Fig. 1.
Toward this aim, one may select for example a linear map $\varPhi(\mathbf{u}_{m})=\mathbf{W}^{\top}\mathbf{u}_{m}$, with $\mathbf{W}\in\mathbb{R}^{N\times d}$ containing the leading Proper Orthogonal Decomposition (POD) modes arising from the covariance matrix of $X$ over all $t_{m}$, yielding the POD projection that maximizes the variance in the $L^{2}$ sense. Nonlinear manifold-learning generalizes this ansatz by constructing $\varPhi$ to preserve the intrinsic geometry of the data: Isomap, for example, builds a neighborhood graph on $\{\mathbf{u}_{m}\}$ and embeds points so that the shortest path distances approximate the manifold geodesics; Diffusion Maps employ Gaussian kernels to define a Markov operator whose leading eigenfunctions embed data such that Euclidean distances in $\mathbb{R}^{d}$ approximate diffusion distances in $\mathcal{M}$; Autoencoder networks learn an encoder $\varPhi_{\theta}$ and a decoder $\Psi_{\phi}:A\rightarrow X$ by minimizing the reconstruction error $\sum_{m}\|\mathbf{u}_{m}-\varPsi_{\phi}(\varPhi_{\theta}(\mathbf{u}_{m}))\|^{2}$, thus capturing complex nonlinear spatio-temporal features. Once $\varPhi$ is obtained, one can learn sparse surrogate models in the latent space via ML. Hence, whether via linear spectral decompositions or flexible nonlinear embeddings, data-driven manifold learning provides a rigorous framework for extracting low-dimensional spatio-temporal structures of fluid flows, the full order models of which are described by the Navier-Stokes equations.
Having the latent variables $\mathbf{a}(t)=\varPhi(\mathbf{u}(t))\in\mathbb{R}^{d}$ obtained at time $t$ via the mapping $\varPhi$, where $\mathbf{u}(t)\in\mathbb{R}^{N}$ does not necessarily correspond to a recorded snapshot $\mathbf{u}_{m}$ used for constructing the mapping, one can learn the governing embedded dynamics using a model in the form:
$\dfrac{d\mathbf{a}(t)}{dt}\;=\;\boldsymbol{\phi}\bigl{(}\mathbf{r}(t),\boldsymbol{\chi}\bigr{)}\;+\;\mathbf{e}(t),$ (3)
where $d\mathbf{a}(t)/dt$ is an estimation of the time-derivative of the latent state, $\mathbf{r}(t)\in\mathbb{R}^{p}$ collects the predictors (e.g., the current and/or past states of the latent dynamics $\mathbf{a}(t),\mathbf{a}(t-\tau_{1}),\dots,\mathbf{a}(t-\tau_{l})$ or other exogenous inputs), $\boldsymbol{\chi}\in\mathbb{R}^{q}$ denotes fixed parameters or known forcings, and $\mathbf{e}(t)$ represents unmodeled time-dependent noise. Learning a surrogate ROM model in the latent space entails selecting a regression map $\mathbf{g}:\mathbb{R}^{p}\times\mathbb{R}^{q}\times\mathbb{R}^{l}\to\mathbb{R}^{d}$ from a hypothesis class $G$ with parameters $\boldsymbol{\theta}\in\mathbb{R}^{l}$, by solving:
$\min_{\mathbf{g}\in G,\;\boldsymbol{\theta}\in\mathbb{R}^{l}}\;\mathcal{L}\!\Bigl{(}\dfrac{d\mathbf{a}_{m}}{dt},\mathbf{g}\bigl{(}\mathbf{r}_{m},\boldsymbol{\chi};\boldsymbol{\theta}\bigr{)}\Bigr{)},\quad\dfrac{d\widehat{\mathbf{a}}_{m}}{dt}=\mathbf{g}\bigl{(}\mathbf{r}_{m},\boldsymbol{\chi};\boldsymbol{\theta}\bigr{)},$ (4)
over the $M$ observations, where $\mathcal{L}$ quantifies the prediction error (e.g., mean-squared error) of the time-derivative estimation. The mapping $\mathbf{g}$ is the second stage of the proposed framework (stage (b) in Fig. 1), which in practice may be chosen from the class of SINDy, neural networks, or GPR models.
The mapping $\varPhi$ to the latent space, and the ROM governing the dynamics therein, enable making predictions, but more importantly performing system-level tasks in low dimensions, which are infeasible for high-dimensions, such as bifurcation analysis. Hereby, we perform numerical continuation to track down stable and unstable stationary states of the fluid flow, on the basis of the constructed ROM; stage (c) in Fig. 1.
Finally, once the surrogate model has produced predictions or stationary states $\hat{\mathbf{a}}_{*}\in\mathbb{R}^{d}$ in the latent space, we need to solve the pre-image (or “lifting”) problem for reconstructing the corresponding high-dimensional velocity field $\hat{\mathbf{u}}_{*}\in\mathbb{R}^{N}$; stage (d) in Fig. 1.
In general, given the forward embedding by the mapping $\varPhi:\mathbb{R}^{N}\to\mathbb{R}^{d}$, we seek a lifting operator $\varPsi$ that maps new latent points $\hat{\mathbf{a}}_{*}\notin\varPhi(X)$ back to the ambient space. This solution to the pre-image (lifting) problem can be formulated as the parameter-dependent minimization:
$\hat{\boldsymbol{c}}\;=\;\arg\min_{\boldsymbol{c}}\Bigl{\|}\,\boldsymbol{a}_{*}\;-\;\varPhi\bigl{(}\varPsi(\boldsymbol{a}_{*};\boldsymbol{c})\bigr{)}\Bigr{\|}_{2},$
subject to any necessary constraints on $\boldsymbol{c}$, where $\varPsi(\cdot\,;\boldsymbol{c}):\mathbb{R}^{d}\to\mathbb{R}^{N}$ is a lifting operator parametrized by $\boldsymbol{c}$. Once $\hat{\boldsymbol{c}}$ is determined, the high-dimensional reconstruction is given by:
$\hat{\boldsymbol{u}}_{*}=\varPsi(\boldsymbol{a}_{*};\,\hat{\boldsymbol{c}}).$ (5)
For a review of several approaches for the solution of the inverse problem, see [43, 44].
For the four-stage data-driven framework described above, we have used POD to find the mapping $\varPhi$ from the high-dimensional space to the latent space, since it offers a closed form solution for the mapping $\varPsi$ of the pre-image problem. To construct the non-intrusive ROM in the latent space, we have implemented GPR models. In the following, we provide details about the data acquisition and discuss each stage of the proposed framework in detail.
2.2 Numerical solution of the Navier-Stokes equations for data acquisition
For acquiring the high-dimensional data of fluid flow upon which the ROM is constructed, we solve the incompressible Navier-Stokes PDEs in Eqs. (1a)-(1c). For each flow configuration, we consider appropriate boundary and initial conditions (see Section 3 for further details) and perform direct numerical simulations by means of the open-source code BASILISK (http://basilisk.fr), which implements a second-order accurate finite-volume scheme. As usual for incompressible flows, Eqs. (1a)-(1c) are solved by means of the so-called projection method, which is here applied on a uniform structured grid of $N_{g}=N_{x}\times N_{y}$ points. In this procedure, a temporary velocity field is first found by ignoring the pressure gradient. In a second step, the temporary field is projected on a space of divergence-free velocity fields by adding the appropriate pressure gradient correction. For a detailed description of the numerical schemes implemented in BASILISK, the reader can refer to [90].
To construct the mapping $\varPhi$ in Eq. (2) across the values of the bifurcation parameter (i.e., in the pre- and post-bifurcation regimes), we perform $N_{Re}$ direct numerical simulations with varying Reynolds number $Re$ values, uniformly distributed in the range $Re\in[Re_{min},Re_{max}]$. The resulting numerical solutions consist of spatio-temporal snapshots of the velocity components $u$ and $v$, collected in the column vectors $\mathbf{u}_{m}\in\mathbb{R}^{2N_{g}}$. For each numerical solution of a fixed $Re$ value, we record $N_{t}$ snapshots of $\mathbf{u}_{m}$ to form the resulting data set, collected in the snapshot matrix $\mathbf{S}=\{\mathbf{u}_{m}\}_{m=1,\ldots,M}\in\mathbb{R}^{N\times M}$, where $N=2N_{g}$ and $M=N_{t}\times N_{Re}$.
2.3 Reduced-order basis construction via Proper Orthogonal Decomposition
As already discussed, for the first stage of the proposed data-driven framework, we find the mapping $\varPhi$ from the high-dimensional space to the latent space in Eq. (2) via manifold learning. Here, for our illustrations, we use POD; as we show, for the particular benchmark problems, it correctly identifies the intrinsic dimension of the normal form of the corresponding bifurcations. For other problems, exhibiting more complex dynamics such as turbulence, we have recently shown [42] that nonlinear manifold learning, such as Diffusion Maps, is more efficient. Hence, here, we describe the methodology based on the POD paradigm.
In particular, for each value of the governing parameter $Re$, one can evaluate the fluctuations of the velocity field $\textbf{u}(x,y,t)=[u(x,y,t),v(x,y,t)]^{\top}$ with respect to the temporal mean $\overline{\textbf{u}}(x,y)$. The celebrated POD technique [91] decomposes such fluctuations as:
$\textbf{u}^{\prime}(x,y,t)=\textbf{u}(x,y,t)-\overline{\textbf{u}}(x,y)=\sum_{i=1}^{\infty}a_{i}(t)\boldsymbol{\varphi}_{i}(x,y),$ (6)
in which the POD modes $\boldsymbol{\varphi}_{i}(x,y)$ are mutually spatially orthogonal. The discrete POD modes $\boldsymbol{\varphi}_{i}$ are obtained by employing the method of snapshots [92], namely by solving the eigenvalue problem formulated for the covariance matrix $\mathbf{S}^{\top}\mathbf{S}\in\mathbb{R}^{M\times M}$, reading as:
$\mathbf{S}^{\top}\mathbf{S}\boldsymbol{\psi}_{i}=\lambda_{i}\boldsymbol{\psi}_{i},\quad i=1,\ldots,M,$ (7)
where $\boldsymbol{\varphi}_{i}=\mathbf{S}\boldsymbol{\psi}_{i}/\sqrt{\lambda_{i}}$ is the $i^{th}$ mode and $\lambda_{i}$, $\boldsymbol{\psi}_{i}$ are the corresponding eigenvalue and eigenvector of the covariance matrix, sorted in a descending order such that $\lambda_{1}>\ldots>\lambda_{M}$. By retaining the leading $d\ll M$ modes, one can construct the reduced order POD basis parameterizing the manifold, which is further used to project the high-dimensional velocity field onto the POD temporal coordinates $a_{i}(t)$, for $i=1,\dots,d$.
Using the POD basis, we define the mapping $\varPhi$ in Eq. (2) as $\mathbf{a}=\varPhi(\mathbf{u})=\mathbf{W}^{\top}\mathbf{u}$, where $\mathbf{W}=[\boldsymbol{\varphi}_{1},\ldots,\boldsymbol{\varphi}_{d}]\in\mathbb{R}^{N\times d}$, thus defining the latent variables as the leading POD coefficients $\mathbf{a}=[a_{1},\ldots,a_{i},\ldots,a_{d}]^{\top}$. Note that the POD basis provides a parameterization of the manifold for seen points $\mathbf{u}_{m}$, but also unseen ones. Hence, it is employed for any $\mathbf{u}(t)$ point in the high-dimensional space, providing the projection in the latent space $\mathbf{a}(t)=\mathbf{W}^{\top}\mathbf{u}(t)$, which will be next used for the ROM construction via GPR models. We hereby highlight that the POD basis provides a closed-form solution for the pre-image mapping $\varPsi$ used for the reconstruction of the high-dimensional space. In particular, for any latent point $\mathbf{a}_{*}$, the high-dimensional reconstruction in Eq. (5) is provided by POD as $\mathbf{u}_{*}=\varPsi(\mathbf{a}_{*})=\mathbf{W}\mathbf{a}_{*}$.
2.4 Gaussian Process regression
Once the reduced coordinates embedding of the latent flow dynamics $\mathbf{a}(t)$ has been obtained, the aim is to identify the unknown evolution operator in the form of Eq. (3). Here, we consider that the predictors $\mathbf{r}(t)$ are simply the latent variables at the current state $\mathbf{a}(t)\in\mathbb{R}^{d}$ and that the fixed parameters $\boldsymbol{\chi}\in\mathbb{R}$ include the bifurcation parameter $Re$. We thus seek to construct a ROM model with inputs $\mathbf{z}(t)=[\mathbf{a}(t),Re]^{\top}\in\mathbb{R}^{d+1}$ and outputs the derivatives $d\mathbf{a}(t)/dt$.
Here, we choose a GPR model from the hypothesis class $G$. Then, the GPR-POD surrogate model for each component of the latent variables $a_{i}$ for $i=1,\ldots,d$ may be approximated as:
$\dfrac{da_{i}}{dt}\;=\;g_{i}\bigl{(}\mathbf{a},Re;\boldsymbol{\theta}_{i}\bigr{)}\;+\;e_{i},\qquad e_{i}\sim\mathcal{N}(0,\sigma_{i}^{2}),$ (8)
where the time derivative $da_{i}/dt$ of the $i$-th latent variable is coupled with all other $\mathbf{a}=[a_{1},\ldots,a_{d}]$; the time-dependence is dropped here for the conciseness of the presentation. Each component $g_{i}(\bm{z};\bm{\theta}_{i})=g_{i}([\mathbf{a},Re]^{\top};\bm{\theta}_{i})$ in Eq. (8), follows an independent (from other components) Gaussian Process $g_{i}\sim\mathcal{G}\mathcal{P}(0,k_{i}(\mathbf{z},\mathbf{z}^{\prime}|\boldsymbol{\theta}_{i}))$ with zero mean and a kernel $k_{i}(\cdot)$ parameterized by $\boldsymbol{\theta}_{i}$ with the inputs $\mathbf{z}$ and $\mathbf{z}^{\prime}$.
For the implementation of GPR, let us consider the POD projections $\mathbf{a}_{m}$ of the $M$ given observations on the latent space, along with the corresponding parameter $Re_{m}$ values, collected in the matrix $\bm{Z}=[\bm{z}_{1},\dots,\bm{z}_{M}]^{\top}\in\mathbb{R}^{M\times(d+1)}$, where $\bm{z}_{m}=[\mathbf{a}_{m},Re_{m}]^{\top}$. Then, the prior distribution of the $i$-th component $g_{i}$ in Eq. (8) is:
$P(\bm{g}_{i}\mid\bm{Z})=\mathcal{N}(\bm{g}_{i}\mid\bm{0},\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})),$ (9)
where $\bm{g}_{i}=[g_{i}(\bm{z}_{1}),\dots,g_{i}(\bm{z}_{M})]^{\top}\in\mathbb{R}^{M}$ collects the $i$-th outputs and $\bm{K}_{i}(\bm{Z},\bm{Z})\in\mathbb{R}^{M\times M}$ is the covariance matrix with entries $k_{i}(\bm{z}_{m},\bm{z}_{l}\mid\bm{\theta}_{i})$ for $m,l=1,\ldots,M$.
Let $d\mathbf{a}_{i}/dt\in\mathbb{R}^{M}$ denote the vector of time derivatives for the $i$-th latent variable, estimated from the $M$ observed snapshots $\{\mathbf{a}_{m}\}_{m=1}^{M}$. Then, predictions at a new point $\bm{z}_{*}\in\mathbb{R}^{d+1}$ are made by drawing $g_{i}(\bm{z}_{*})$ from the joint distribution:
$\begin{bmatrix}\frac{d\mathbf{a}_{i}}{dt}\\ g_{i}(\bm{z}_{*})\end{bmatrix}\sim\mathcal{N}\left(\bm{0},\begin{bmatrix}\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})+(\sigma_{i})^{2}\bm{I}_{M}&\bm{k}_{i}(\bm{Z},\bm{z}_{*}\mid\bm{\theta}_{i})\\ \bm{k}_{i}(\bm{z}_{*},\bm{Z}\mid\bm{\theta}_{i})&k_{i}(\bm{z}_{*},\bm{z}_{*}\mid\bm{\theta}_{i})\end{bmatrix}\right).$ (10)
It can be shown that the posterior conditional distribution $P(g_{i}(\bm{z}_{*})|d\mathbf{a}_{i}/dt,\bm{Z},\bm{z}_{*})$ follows a normal distribution $g_{i}(\bm{z}_{*})\sim\mathcal{N}(\mu_{i,*},\sigma_{i,*}^{2})$ with the expected value and covariance of the estimation given by:
$\mu_{i,*}$ $=\bm{k}_{i}(\bm{z}_{*},\bm{Z}\mid\bm{\theta}_{i})\left[\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})+\sigma_{i}^{2}\bm{I}_{M}\right]^{-1}\tfrac{d\mathbf{a}_{i}}{dt},$ (11)
$\sigma_{i,*}^{2}$ $=k_{i}(\bm{z}_{*},\bm{z}_{*}\mid\bm{\theta}_{i})-\bm{k}_{i}(\bm{z}_{*},\bm{Z}\mid\bm{\theta}_{i})\left[\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})+\sigma_{i}^{2}\bm{I}_{M}\right]^{-1}\bm{k}_{i}(\bm{Z},\bm{z}_{*}\mid\bm{\theta}_{i}).$ (12)
The hyperparameters $\bm{\theta}_{i}$ and noise variance $\sigma_{i}^{2}$ in the above equations are estimated by minimizing the negative log marginal likelihood (NLML):
$-\log P\left(\tfrac{d\mathbf{a}_{i}}{dt}\mid\bm{Z},\bm{\theta}_{i}\right)=\frac{1}{2}\left(\tfrac{d\mathbf{a}_{i}}{dt}\right)^{\top}\bm{\Sigma}_{i}^{-1}\tfrac{d\mathbf{a}_{i}}{dt}+\frac{1}{2}\log|\bm{\Sigma}_{i}|+\frac{M}{2}\log 2\pi,$ (13)
where $\bm{\Sigma}_{i}=\bm{K}_{i}(\bm{Z},\bm{Z}\mid\bm{\theta}_{i})+\sigma_{i}^{2}\bm{I}_{M}$. For our analysis, we have employed a radial basis function kernel with automatic relevance determination (ARD):
$k_{i}(\bm{z}_{m},\bm{z}_{l}\mid\bm{\theta}_{i})=(\theta_{i,1})^{2}\exp\left(-\sum_{j=1}^{d+1}\frac{(z_{m,j}-z_{l,j})^{2}}{2(\theta_{i,j+1})^{2}}\right),$ (14)
where $z_{m,j}=a_{m,j}$ for $j=1,\ldots,d$ and $z_{m,d+1}=Re_{m}$, and the hyperparameters $\bm{\theta}_{i}=[\theta_{i,1},\dots,\theta_{i,d+2}]$ govern the output scale and the input sensitivities.
Finally, considering the POD projections $\mathbf{a}_{m}$ and the respective $Re_{m}$ values, we first estimate the time derivatives $da_{m,i}/dt$ using finite differences. Then, we determine the parameters $\bm{\theta}_{i}$ of each $g_{i}$ in Eq. (8) by minimizing the NLML in Eq. (13). The resulting POD-GPR model for the dynamics in the latent space takes the form:
[…]
2.5 Numerical bifurcation analysis
The identification of the POD-GPR model of the latent space dynamics in Eq. (8) enables tracking the stationary states and identifying their stability for the construction of the coarse-grained bifurcation diagram.
Typically, the stable stationary states are obtained by numerically finding the roots $\mathbf{a}_{*}$ of:
$\mathbf{g}(\mathbf{a},Re)=\mathbf{0},$ (16)
using a Newton-Raphson scheme, where $\mathbf{g}=[g_{1},\ldots,g_{d}]$ is the learned latent dynamics operator. For tracking unstable stationary states, one needs to continue branches of varying $Re$ parameter values. This can be achieved with pseudo arc-length continuation, by augmenting Eq. (16) with the condition:
$N(\mathbf{a},Re)=\mathbf{c}\cdot\left(\mathbf{a}-\mathbf{a}_{1}\right)+b(Re-Re_{1})=0,\qquad\mathbf{c}\equiv\dfrac{\left(\mathbf{a}_{1}-\mathbf{a}_{0}\right)^{\top}}{\delta s},\qquad b\equiv\dfrac{Re_{1}-Re_{0}}{\delta s}$ (17)
where $(\mathbf{a}_{0},Re_{0})$ and $(\mathbf{a}_{1},Re_{1})$ are two previous computed stationary states (satisfying Eq. (16)) along the branch and $\delta s$ is the pseudo arc-length step. The condition in Eq. (17) constrains the next stationary state along the branch of Eq. (16) to lie on a hyperplane orthogonal to the tangent of the branch at $(\mathbf{a}_{1},Re_{1})$; the tangent is approximated using the previous point along the branch $(\mathbf{a}_{0},Re_{0})$.
In practice, the next stationary state is obtained by numerically finding the roots of Eqs. (16)-(17) via the iterative solution of the linearized $(d+1)$-dimensional system:
$\begin{bmatrix}\dfrac{\partial\mathbf{g}}{\partial\mathbf{a}}&\dfrac{\partial\mathbf{g}}{\partial Re}\\ \mathbf{c}&b\end{bmatrix}\begin{bmatrix}\delta\mathbf{a}\\ \delta Re\end{bmatrix}=-\begin{bmatrix}\mathbf{g}(\mathbf{a},Re)\\ N(\mathbf{a},Re)\end{bmatrix},$ (18)
where $\partial\mathbf{g}/\partial\mathbf{a}\in\mathbb{R}^{d\times d}$ and $\partial\mathbf{g}/\partial Re\in\mathbb{R}$ are the Jacobian matrix and the derivatives of the stationary state condition in Eq. (16), respectively. At each iteration, the system in Eq. (18) is solved with Moore-Penrose pseudo-inverse for computing the updates $\delta\mathbf{a}$ and $\delta Re$, until reaching a predefined tolerance criterion for $\mathbf{g}(\mathbf{a},Re)$ and $\delta\mathbf{a}$. The resulting converged solution $(\mathbf{a}_{*},Re_{*})$ is then used for taking the next step along the stationary state branch; i.e., by setting $(\mathbf{a}_{1},Re_{1})=(\mathbf{a}_{0},Re_{0})$ and $(\mathbf{a}_{*},Re_{*})=(\mathbf{a}_{1},Re_{1})$ in Eq. (18).
Once obtaining the stationary state $(\mathbf{a}_{*},Re_{*})$, one can determine its stability by the eigenvalues $\sigma_{i}$ for $i=1,\ldots,d$ of the Jacobian matrix $\partial\mathbf{g}/\partial\mathbf{a}$. When the $\operatorname{Re}(\sigma_{i})<0$ for all $i=1,\ldots,d$, the stationary state is stable, while when $\operatorname{Re}(\sigma_{i})>0$ for some $i$, the stationary state is unstable. In the case where $\operatorname{Re}(\sigma_{i})=0$, a bifurcation point is located along the equilibrium branch. Codimension-1 bifurcations (when one eigenvalue crosses the imaginary axis), such as Hopf, fold, or branching points, mark critical transitions where stability changes and new branches emerge. Hopf bifurcations give rise to branches of limit cycles (isolated periodic orbits), while branching points give rise to additional equilibrium branches. Detailed discussion about codimension-2 bifurcations (encountered in biparametric bifurcation analysis) and the criteria for locating bifurcation points can be found in [78].
If a branch of limit cycle emerges, one seeks periodic solutions satisfying $\mathbf{a}(t)=\mathbf{a}(t+T)$, $\forall t$ where $T$ is the period of the orbit. In this case, $T$ is treated as a parameter and the system of the latent space dynamics in Eq. (8) is rescaled in time $\tau=t/T$, so that the rescaled period is $1$. Then, one has to solve the following boundary value problem:
[…]
where the unknown period $T$ has to be determined, along with the limit cycle state $\mathbf{a}_{*}$. To avoid translational invariance over the infinite periodic solutions, the above system is augmented by a phase condition:
[…]
where $\dot{\mathbf{a}}_{0}$ is the derivative of a previously computed limit cycle. Eqs. (19)-(20) together form a system $\mathbf{G}(\mathbf{a},Re,T)$ in the form of Eq. (16), the solution of which determines the limit cycle orbit, the bifurcation parameter and the corresponding period. Continuation of the limit cycle branch is performed using a similar procedure as the one followed for the stationary states, only in the former case one has to first discretize $\mathbf{a}$ over a set of collocation points. Finally, once obtaining a limit cycle orbit $(\mathbf{a}_{*},Re_{*},T_{*})$, its stability is determined by the Floquet multipliers, which also provide information for the identification of bifurcation points along the limit cycle branch, including period doubling, fold, branching or Neimark-Sacker bifurcation points. For a detailed presentation of the limit cycle continuation and bifurcation, see [78].
For the implementation of the numerical bifurcation analysis to the POD-GPR model, we have employed the software package MATCONT [3, 4], which provides a complete toolkit for continuation and identification of bifurcation points, both for stationary states and limit cycles. In particular, we supplied the learned POD-GPR model in Section 2.4, computing numerically the derivatives required in Eq. (18).
We highlight that MATCONT employs direct numerical solvers for solving the system in Eq. (18), making numerical bifurcation analysis of the full Navier-Stokes equations computationally intractable, particularly for limit cycles. While matrix-free Krylov methods [5, 8, 9, 10] can solve Eq. (18) for large systems, they require custom implementations that lack the comprehensive bifurcation analysis capabilities of specialized packages like MATCONT. Our proposed framework addresses this limitation by performing all continuation and bifurcation analysis in the latent space through the POD-GPR model, and then reconstructing solutions to the high-dimensional space via the mapping $\varPsi$ described in Section 2.3.
3 Flow configurations
For our illustration, we focus on two different flow configurations: the wake flow past an infinitely long circular cylinder and the sudden-expansion channel flow. The physical layouts with the associated flow fields and regimes are described in this section.
3.1 Wake flow past a cylinder
The two-dimensional flow arising around an infinitely long circular cylinder immersed in a uniform stream is schematically represented in Fig. 2(a). A Cartesian coordinate system $\mathcal{O}x^{\star}y^{\star}$ is placed in the cylinder centre, with the $x^{\star}$ axis pointing in the flow direction, the $y^{\star}$-axis along the normal-to-flow direction, and the $z^{\star}$-axis (not represented) running along the cylinder centreline. The cylinder diameter is denoted as $D^{\star}$, while the free-stream velocity (aligned with the $x^{\star}$ direction) is $U^{\star}_{\infty}$. The Reynolds number for this setup is defined as
$Re=\dfrac{\rho U^{\star}_{\infty}D^{\star}}{\mu},$ (21)
where $\rho$ and $\mu$ are the density and dynamic viscosity of the fluid, respectively. Note that all dimensional quantities, except the fluid properties ($\rho,\mu$), are denoted with the superscript $\star$.
Figure 2: Schematic representations of the flow past an infinitely long circular cylinder (panel (a)) and the planar sudden-expansion channel flow (panel (b)).
Figure 3: Instantaneous visualization of $u$ (panels (a) and (c)) and $v$ (panels (b) and (d)) velocity components contour maps of the cylinder flow for $Re=20$ ((a)-(b)) and $Re=60$ ((c)-(d)). Panel (e) reports the instantaneous spatial distribution of $u$ on the $y=0$ axis for $Re=20$ (red curve), $Re=40$ (blue curve), and $Re=60$ (green curve). Panel (f) shows the temporal evolution of $u$ at the streamwise location $(\tilde{x},\tilde{y})=(7,0)$ for the same values of Reynolds number.
Direct numerical simulations of the incompressible Navier–Stokes equations are used to compute the two-dimensional viscous wake behind the cylinder (see previous Section 2.2 for numerical implementation details). All physical quantities are made dimensionless with respect to the diameter $D^{\star}$ and velocity $U^{\star}_{\infty}$,
$x=\dfrac{x^{\star}}{D^{\star}},\quad y=\dfrac{y^{\star}}{D^{\star}},\quad u=\dfrac{u^{\star}}{U^{\star}_{\infty}},\quad v=\dfrac{v^{\star}}{U^{\star}_{\infty}},\quad p=\dfrac{2p^{\star}}{\rho U^{\star 2}_{\infty}}\quad t=t^{\star}\dfrac{U^{\star}_{\infty}}{D^{\star}}.$ (22)
As shown in Fig. 2(a), the computational domain is a rectangle excluding the interior of the cylinder, with sides equal to $L^{\star}_{x}=26D^{\star}$ and $L^{\star}_{y}=12D^{\star}$. As the cylinder is considered fixed in the present analysis, a no-slip boundary condition is applied on its surface. At the domain inlet (left side of the rectangle), a uniform velocity profile is prescribed, namely $u=1$ and $v=0$, where $u$ and $v$ are the streamwise and normal-to-flow velocity components, respectively. The same values are also assigned as initial conditions to start the computations. A standard free-outflow boundary condition is enforced on the domain outlet (right side of the rectangle), while the remaining sides are equipped with homogeneous Neumann boundary conditions for all variables. A uniform structured grid is employed to discretize the physical domain, with mesh size equal to $\Delta x=\Delta y=0.2$. This corresponds to 5 grid cells within the length scale $D^{\star}$, namely $N_{g}=N_{x}\times N_{y}=7800$ total number of grid points. Note that such a spatial resolution has been shown to be adequate to reproduce the 2D cylinder wake flow dynamics [93].
Spatio-temporal numerical simulations are performed in the range $Re\in[20,70]$. The streamwise $u$ and normal-to-flow $v$ velocity components are stored with a time-step $\Delta t=0.2$, for a total computational time equal to $T=400$. Therefore, 2000 temporal realizations of the velocity field are considered for each value of the Reynolds number.
Snapshots of the simulations are shown in Fig. 3 in terms of instantaneous two-dimensional contour maps of the streamwise (panels (a) and (c)) and normal-to-flow (panels (b) and (d)) velocity components by varying the Reynolds number: $Re=20$ (panels (a)-(b)) and $Re=60$ ((c)-(d)). Moreover, Fig. 3(e) reports the spatial distribution $u(x)$ along the symmetry axis $y=0$ corresponding to $Re=20$, $40$ and $60$, while Fig. 3(f) shows the temporal evolution of $u$ at the streamwise location $(\tilde{x},\tilde{y})=(7,0)$, for the same $Re$ values.
For relatively low Reynolds number ($Re=20$ and $Re=40$), the flow is in a steady regime, and it is characterized by a recirculation wake region, i.e. a flow area where $u<0$. As one can appreciate by comparing red and blue curves in Fig. 3(e), the extension of the wake increases by increasing $Re$. As the Reynolds number increases up to $Re=60$, the velocity field exhibits periodic spatio-temporal oscillations due to the vortex shedding phenomenon originating behind the cylinder (see Fig. 3(c)-(d) and green curves in Fig. 3(e)-(f)). The onset of the periodic regime is determined by the instabilities developed in the recirculation wake region near $Re_{cr}\approx 49$, which are known to be the manifestation of an Andronov-Hopf bifurcation [94].
3.2 Sudden-expansion channel flow
Figure 4: Instantaneous visualization of $u$ (panels (a) and (c)) and $v$ (panels (b) and (d)) velocity components contour maps of the sudden-expansion channel for $Re=30$ ((a)-(b)) and $Re=60$ ((c)-(d)). The white curves in panels (a)-(d) denote the flow locations where $u=0$, namely the wake borders. Panels (e)-(f) report the time variation of the upper $x^{+}_{w}$ (black curve) and lower $x^{-}_{w}$ (red curve) wake regions extension for $Re=30$ and $Re=60$, respectively.
The second flow configuration considered in this work is sketched in Fig. 2(b). It consists of a fluid stream developing in a planar two-dimensional channel, which undergoes a sudden expansion of the cross section as the streamwise direction $x^{\star}$ increases. The channel has length $L^{\star}_{c}$ and the coordinate system $\mathcal{O}x^{\star}y^{\star}$ is placed on the symmetry axis at the inlet section. The sudden expansion of the cross-sectional unit area from $A^{\star}=h^{\star}$ to $A^{\star}=H^{\star}$ is located at $x^{\star}=L^{\star}_{e}$. As in previous works [95, 96], the expansion ratio $W=H^{\star}/h^{\star}$ is set equal to $W=6$, and the channel entrance $L^{\star}_{e}$ and total $L^{\star}_{c}$ lengths are assigned equal to $L^{\star}_{e}=12h^{\star}$ and $L^{\star}_{c}=60h^{\star}$, respectively. These values ensure that the fluid stream development in the entrance region does not affect the flow regimes established in the downstream part of the channel. The Reynolds number for this setup is defined as
$Re=\dfrac{2\rho U^{\star}_{e}h^{\star}}{\mu},$ (23)
where $U^{\star}_{e}$ is the averaged velocity at the inlet section, i.e. at $x^{\star}=0$.
As for the cylinder wake flow configuration, we compute the flow regimes characterizing the sudden-expansion channel flow by direct numerical simulations (implementation details are reported in the previous Section 2.2). All physical quantities are made dimensionless with respect to the entrance channel height $h^{\star}$ and averaged inlet velocity $U^{\star}_{e}$,
$x=\dfrac{x^{\star}}{h^{\star}},\quad y=\dfrac{y^{\star}}{h^{\star}},\quad u=\dfrac{u^{\star}}{U^{\star}_{e}},\quad v=\dfrac{v^{\star}}{U^{\star}_{e}},\quad p=\dfrac{2p^{\star}}{\rho U^{\star 2}_{e}}\quad t=t^{\star}\dfrac{U^{\star}_{e}}{h^{\star}}.$ (24)
At the domain inlet (left side of the entrance channel), a parabolic velocity profile with mean value equal to $U^{\star}_{e}$ is prescribed. A standard free-outflow boundary condition is enforced on the domain outlet (right side of the channel), while the remaining sides are equipped with no-slip boundary conditions, namely $u=v=0$. The same values are also prescribed as initial conditions to start the computation. A uniform, structured grid is employed to discretize the physical domain, with mesh size equal to $\Delta x=\Delta y=0.12$. This corresponds to 9 grid cells within the length scale $h^{\star}$, namely $N_{g}=N_{x}\times N_{y}=26214$ total number of grid points, which was necessary to achieve the grid-independence of the flow regimes computed by variation of the Reynolds number $Re$.
Spatio-temporal numerical simulations are performed in the range $Re\in[30,70]$. The streamwise $u$ and normal-to-flow $v$ velocity components are stored with a time-step $\Delta t=0.5$, for a total computational time equal to $T=1600$. Therefore, 3200 temporal realizations of the velocity field are considered for each value of the Reynolds number.
Figure 5: Symmetric unstable (panel (a)) and asymmetric conjugated stable (panels (b)-(c)) solutions of the sudden-expansion channel flow in terms of streamwise velocity $u$ contour maps. The white curves in all panels denote the flow locations where $u=0$, namely the wake borders. Here, $Re=70$.
Snapshots of the simulations are shown in Fig. 4 in terms of instantaneous two-dimensional contour maps of the streamwise (panels (a) and (c)) and normal-to-flow (panels (b) and (d)) velocity components by varying the Reynolds number: $Re=30$ (panels (a)-(b)) and $Re=60$ ((c)-(d)).
For the lowest Reynolds number considered, a steady symmetric flow with respect to the $y=0$ direction is observed, which is characterized by two recirculation (wake) regions of equal size developing downstream of the expansion (Fig. 4(a)-(b)). As the Reynolds number increases, the flow symmetry is initially maintained, and the wake regions progressively increase. This aspect is quantified in Fig. 4(e), which reports the upper $x^{+}_{w}$ and lower $x^{-}_{w}$ wake regions extension for $Re=30$ as the computational time $t$ increases. Above a certain critical threshold denoted here as $Re_{sb}$, the symmetry of the velocity field breaks due to the so-called Coanda effect [97]: a localized increase in velocity near one wall determines a reduction of the local pressure, which establishes a pressure difference maintained across the channel. As a consequence, one of the downstream recirculation regions expands and the other shrinks, resulting in the steady asymmetric solution observed at $Re=60$ (see Fig. 4(c)-(d) and Fig. 4(f)).
The symmetry breaking occurs as a result of a supercritical pitchfork bifurcation in the solution of the Navier–Stokes equations [98]. In other terms, above $Re_{sb}$ two asymmetric conjugated stable solutions co-exist, while the symmetric solution becomes unstable (see Fig. 5). Note that the threshold $Re_{sb}$ has been found to be dependent on the channel expansion ratio $W$ [95, 96]. For the geometry considered here ($W=6$) and the Reynolds number definition given by Eq. (23), the value is $Re_{sb}\approx 44$.
4 Results
The POD-GPR methodology introduced in Section 2 is here applied to learn the surrogate reduced-order models of the latent dynamics of the cylinder and channel flow configurations (Sections 4.1-4.2). The ROMs are then employed to perform bifurcation and stability analysis tasks in Sections 4.3-4.4.
4.1 Reduced-order basis identification
Figure 6: Eigenvalues spectrum and corresponding cumulative sum of the cylinder flow (panels (a)-(b), respectively). Panels (c)-(d), (e)-(f) and (g)-(h) report the leading spatial modes $\varphi_{1}$, $\varphi_{2}$ and $\varphi_{3}$ for $u$ (left panels) and $v$ (right panels) fields, respectively. Panels (i)-(j) show the temporal evolution of the velocity field projections $a_{i}$ on the five leading modes $\varphi_{i}$ ($i=1,\dots,5$) for $Re=40$ and $60$, respectively.
Figure 7: Eigenvalues spectrum and corresponding cumulative sum of the sudden-expansion channel flow (panels (a)-(b), respectively). Panels (c)-(d), (e)-(f) and (g)-(h) report the leading spatial modes $\varphi_{1}$, $\varphi_{2}$ and $\varphi_{3}$ for $u$ (left panels) and $v$ (right panels) fields, respectively.
Figure 8: Temporal evolution of the sudden-expansion channel flow velocity field projections $a_{i}$ on the five leading modes $\varphi_{i}$ ($i=1,\dots,5$) for $Re=30$ (panel (a)) and $Re=70$: upward-deflected (panel (b)) and downward-deflected (panel (c)) solutions.
The POD eigenvalues spectrum and the corresponding cumulative sum are represented in Fig. 6(a)-(b) and Fig. 7(a)-(b), respectively for the cylinder and the channel flows. Moreover, the three leading modes $\varphi_{1}$, $\varphi_{2}$ and $\varphi_{3}$ of the two configurations are displayed in panels (c)-(d), (e)-(f), and (g)-(h), respectively, where left panels refer to $u$ and right panels to $v$ velocity fields. The projection of the high-dimensional flow field on the five leading modes are shown in Fig. 6(i)-(j) and Fig. 7(i)-(j), for Reynolds number values both below (left panels) and above (right panels) the critical thresholds $Re_{cr}\approx 49$ and $Re_{sb}\approx 44$. Values of the parameters employed to build the POD snapshot matrix (see previous Section 2.3) corresponding to the cylinder flow are $Re_{min}=40$, $Re_{max}=60$, $N_{g}=7800$, $N_{t}=1000$, $N_{Re}=11$, for a total number of snapshots equal to $M=22000$. For the channel flow configuration, we have employed $Re_{min}=30$, $Re_{max}=70$, $N_{g}=26214$, $N_{t}=3200$, $N_{Re}=21$, for a total number of snapshots equal to $M=134400$. flow simulations wih the cylinder setup, it can be seen that the first two modes $\boldsymbol{\varphi_{1}}$ (Fig. 6(c)-(d)) and $\boldsymbol{\varphi_{2}}$ (Fig. 6(e)-(f)) are physically related to the periodic von Karman street of vortices establishing downstream of the cylinder for $Re>Re_{cr}$. The third mode $\boldsymbol{\varphi_{3}}$ (Fig. 6(g)-(h)) is associated with a shift-mode characteristic of the transient dynamics from the onset of vortex shedding to the periodic von Karman wake [99]. The fourth $\boldsymbol{\varphi_{4}}$ and fifth $\boldsymbol{\varphi_{5}}$ modes (not shown) represent higher harmonics of the leading two modes, and thus they do not describe new directions along the data set. By looking at the temporal evolution of the velocity field projections $a_{i}(t)$ for $Re=40$ (Fig. 6(i)) and $Re=60$ (Fig. 6(j)), it can be clearly appreciated that the leading two modes $\boldsymbol{\varphi_{1}}$ and $\boldsymbol{\varphi_{2}}$ are sufficient to describe the bifurcation phenomenon separating the steady and the periodic flow regimes. For this reason, out of the total number of modes, we retain the first $P=2$ to learn the surrogate Gaussian Process regression model suitable for the bifurcation analysis of the cylinder flow dynamics. This is in line with the dimension of the normal form for a Hopf–Andronov bifurcation.
For the numerical simulations on the suddenly-expanding channel, one can easily observe that the velocity field projection of the first mode $\boldsymbol{\varphi_{1}}$ (Fig. 7(c)-(d) and black curves in Fig. 8(a)-(c)) embeds the symmetry-breaking bifurcation experienced by the flow as the Reynolds number increases. The temporal evolution of the first latent-dynamics coordinate indeed converges towards a unique fixed-point solution $a_{1}=0$ for $Re=30$ (Fig. 8(a)), and to mirrored solutions $a_{1}=\pm 16.5$ for $Re=70$ (Fig. 8(b)-(c)), namely below and above the threshold $Re_{sb}\approx 44$, respectively. On the other hand, the remaining coordinates $a_{i}$ ($i=2,\dots,5$) exhibit in the long time limit approximately the same values for the two conjugated solutions existing for $Re>Re_{sb}$ (see red, blue, magenta and green curves in Fig. 8(b)-(c)). Therefore, only the first $P=1$ mode will be considered to learn the surrogate model of the sudden-expansion channel flow dynamics. In that way, again this is in agreement with the dimension of the norm-form of the pitchfork bifurcation that appears in this range of the bifurcation parameter.
4.2 Gaussian Process ROMs
Figure 9: Numerical data set for learning the GPR model of the cylinder flow latent dynamics.
Figure 10: Numerical data set for learning the GPR model of the sudden-expansion channel flow latent dynamics.
[…]
Figure 11: Accuracy of the GPR model on the testing data set $d\textbf{a}/dt$ for the cylinder (panel (a)) and channel (panel (b)) flow configurations. For given input data $(Re,\textbf{a})\equiv(Re,a_{1},\dots,a_{P})$, the model predictions are shown both in terms of expected (mean) values (black dots) and 95% CI (confidence intervals, cyan band). The black dashed line in both panels denotes the quadrant bisector.
4.3 Andronov-Hopf bifurcation
The numerical integration of the reduced-order model of the cylinder flow dynamics is first performed for two different values of the Reynolds number, spanning both steady laminar ($Re<Re_{cr}$) and periodic vortex shedding ($Re>Re_{cr}$) regimes. Results are shown in Fig. 12, where the time-evolution of the reduced coordinates $a_{1}$ and $a_{2}$ in the phase-space is reported for $Re=45$ (panel (a)) and $Re=55$ (panel (b)). Note that values from the test data set, i.e., unseen during the training step, have been employed for the Reynolds number and for the initial conditions $a_{1}(t=0)$ and $a_{2}(t=0)$.
Figure 12: Reduced-order dynamics of the cylinder wake flow as predicted by time integration of the GPR model (15) by variation of the Reynolds number: $Re=45$ (panel (a)); $Re=55$ (panel (b)).
From the analysis of Fig. 12, it can be appreciated that the learned surrogate model qualitatively reproduces the cylinder wake flow dynamics, both in steady and periodic regimes. The temporal evolutions of $a_{1}$ and $a_{2}$ reduced coordinates asymptotically converge to the fixed-point stable solution $a_{1}=a_{2}=0$ for $Re=45$ (Fig. 12(a)), while limit cycle periodic oscillations are recovered for $Re=55$ (Fig. 12(b)).
[…]
Once the performance of the GPR reduced-order model has been validated through predictions of flow dynamics under new conditions, a numerical bifurcation diagram is constructed using the continuation method described in Section 2.5. The results are presented in Fig. 13, where the onset of a Hopf bifurcation is clearly identified at the critical Reynolds number $Re_{cr}=48.48$.
For $Re<Re_{cr}$, the system exhibits a single stable solution—a fixed-point attractor—represented by the overlapping solid black and red curves $a_{1}=a_{2}=0$ in Fig. 13(a). As the Reynolds number increases beyond $Re_{cr}$, this fixed point becomes unstable to small perturbations, as indicated by the dashed curves in Fig. 13(a). Therefore, a branch of stable limit cycles (LC) emerges from the Hopf bifurcation point, shown as a blue curve in the same panel. In Fig. 13(c), the eigenvalues of the Jacobian matrix associated with system (15) are reported as a function of $Re$. As characteristic of a Hopf bifurcation, a complex-conjugate pair of eigenvalues (red curves) crosses the imaginary axis and becomes unstable (i.e., the real part $\mathrm{Re}(\sigma_{i})>0$ for $i=1,2$, black curve) for $Re>Re_{cr}$. Note also that the imaginary part of the eigenvalues, which corresponds to the oscillation frequency of the resulting limit cycles, increases slightly with $Re$ in the range considered here. This trend is consistent with previous findings in the literature [93].
The continuation along the limit cycles branch enables the computation of the oscillation period $T$ (black curve in Fig. 13(d)) and the associated Floquet multipliers $f_{1}$ and $f_{2}$ (blue curves in Fig. 13(d)), which characterize the stability of periodic solutions. The first Floquet multiplier $f_{1}$ is equal to one at the bifurcation point and decreases monotonically with increasing $Re$. This behavior reflects the increasing stability of the limit cycle as the Reynolds number grows, since a smaller magnitude of $f_{1}$ indicates faster decay of perturbations transverse to the periodic orbit. In contrast, the second Floquet multiplier $f_{2}$ remains constant and equal to 1 throughout the branch. This is an expected trivial result, since $f_{2}=1$ corresponds to the neutral direction associated with the time invariance of the periodic orbit, a characteristic property of limit cycles. The values $f_{1}=f_{2}=1$ at the bifurcation point are also consistent with the theoretical prediction for a supercritical Hopf bifurcation.
Finally, by addressing the pre-image problem, the POD-GPR reduced-order solutions can be lifted back to the original high-dimensional physical space to reconstruct the full bifurcation diagram, which is shown in Fig. 13(b). For comparison, we have also reported the Navier-Stokes solutions (black dots in Fig. 13(b)) in terms of transverse velocity $v$ stored at the spatial location $x=10$, $y=0$. Note that for $Re>Re_{cr}$, both the maximum and minimum amplitude of $v$ oscillations have been considered at the selected spatial location. It can be seen that the high-dimensional bifurcation diagram is accurately reconstructed by the minimal $P=2$ reduced-order model in the whole Reynolds number range considered, with a maximum relative percentage error equal to 12% for $Re=50$.
4.4 Symmetry-breaking bifurcation
Figure 14: Panel (a): reduced-order dynamics of the sudden-expansion channel flow as predicted by time integration of the GPR model (15) by variation of the Reynolds number: $Re=35$ (dashed curves); $Re=55$ (continuous curves). Note that black and red curves refer to different initial conditions $a_{1}(t=0)$. Panels (b)-(c): bifurcation diagram of the sudden-expansion channel flow reduced-order model by increasing the Reynolds number $Re$. In panel (b), the computation is directly performed on the Gaussian Process regression model (15), while in panel (c) the transformation (25) is enforced before computing the solution branches. In panel (d), the high-dimensional bifurcation diagram corresponding to $P=1$ (black continuous curve) and $P=3$ (blue continuous curve) reduced-order coordinates is reconstructed in terms of $v$ velocity component at the spatial location $x=20$, $y=0$. Values from the Navier-Stokes solutions are also reported for comparison (black dots).
The numerical integration of the reduced-order model of the sudden-expansion channel flow dynamics is first performed for two different values of the Reynolds number, spanning the pre- ($Re<Re_{sb}$) and post-bifurcation ($Re>Re_{sb}$) regimes. The results are shown in Fig. 14(a), where the time evolution of the reduced coordinate $a_{1}$ is reported for $Re=35$ (dashed curves) and $Re=55$ (solid curves). Note that values from the test data set, that is,is, unseen during the training step, have been used for the Reynolds number, as well as for the initial condition $a^{0}_{1}\equiv a_{1}(t=0)$. Moreover, note that two conjugated initial conditions have been considered for each $Re$, namely $a^{0}_{1}=-2$ (black curves) and $a^{0}_{1}=2$ (red curves).
From the analysis of Fig. 14(a), it can be seen that the learned surrogate model qualitatively reproduces the sudden-expansion channel flow dynamics, both in the pre- and post-bifurcation regimes. Regardless of the assigned initial condition, the temporal evolution of the reduced coordinate $a_{1}$ converges to the (only) fixed-point stable solution $a_{1}=0$ for $Re=35$ (dashed curves in Fig. 14(a)), after a transient time equal to $t\approx 400$. On the other hand, depending on the value of the initial condition $a^{0}_{1}$, the upward ($a_{1}<0$) and downward ($a_{1}>0$) deflected solutions are asymptotically recovered for $Re=55$ (solid curves in Fig. 14(a)).
Once the performances of the reduced-order GPR model have been assessed by forecasting the flow dynamics for new conditions, the numerical bifurcation diagram is first computed by applying to the system (15) the continuation method described in Section 2.5. As a result, we obtain the bifurcation diagram shown in Fig. 14(b), which identifies the existence of three steady solutions — two stable (continuous red curves) and one unstable (dashed red curve) — for $Re>43.12$. For $Re<43.12$, a unique stable steady solution exists. However, the flip symmetry of the generic pitchfork bifurcation is clearly retrieved in the computed diagram. Due to the intrinsic systematic error of the learned model, we have indeed identified a perturbed pitchfork, in which the lower branch remains permanently stable, and the upper branch exhibits a turning point where a stable and an unstable solution meet. As recently shown by [100] in the context of residual neural network models of agent-based systems, a simple odd symmetry transformation of Eq. (15) is sufficient to recover the pitchfork bifurcation. In particular, given the identified right-hand side $g(Re,a_{1})$ of the Gaussian Process regression model, we compute
$\overline{g}(Re,a_{1})=\dfrac{g(Re,a_{1})-g(Re,-a_{1})}{2}.$ (25)
The transformed model (25) is then used to construct the bifurcation diagram shown in Fig.14(c), which correctly identifies the symmetry-breaking pitchfork bifurcation at $Re_{sb}=43.12$.
Finally, by addressing the pre-image problem, the POD-GPR reduced-order solutions can be lifted back to the original high-dimensional physical space to reconstruct the full bifurcation diagram, which is shown in Fig. 14(d). For comparison, we have also reported the Navier-Stokes solutions (black dots in Fig. 14(d)) in terms of transverse velocity $v$ stored at the spatial location $x=20$, $y=0$. It can be seen that the high-dimensional bifurcation diagram is already well reconstructed by the minimal $P=1$ reduced-order model (black curve in Fig. 14(d)) in the whole Reynolds number range considered, with a maximum relative percentage error equal to 22% for $Re=70$. The accuracy of the reconstruction improves by reducing the maximum error down to 3% if one takes into account the leading $P=3$ reduced-order coordinates in the surrogate POD-GPR model (blue curve in Fig. 14(d)).
5 Conclusions
Here, we presented a four-stage, fully data-driven framework using manifold learning and GPR to construct surrogate (non-intrusive) “normal-form” ROMs with minimal dimensionality and structure, enabling accurate approximation of emergent Navier–Stokes dynamics for bifurcation and stability analyses. In the first step, we have applied manifold learning to extract a low-dimensional representation of the complex, high-dimensional flow fields over the parameter space generated by Navier–Stokes simulations. In the second stage, the temporal evolution of the latent dynamics was learned by using GPR models, allowing also for uncertainty quantification. Then, based on the constructed reduced-order models, the toolkit of numerical bifurcation analysis theory has been exploited to construct the bifurcation diagrams in the low-dimensional latent spaces. Such an analysis is, for all practical purposes, infeasible for high-dimensional systems such as the Navier-Stokes equations. Finally, by solving the pre-image problem, we reconstructed accurately the stationary states of the flow in the original high-dimensional space. For our illustration, the proposed methodology has been demonstrated by focusing on two suitable test-bed flow configurations: the wake flow past an infinitely long circular cylinder and the planar sudden-expansion channel flow. The proposed methodology was shown to achieve high accuracy in reconstructing the bifurcation diagram for such benchmark problems.
The potential of the proposed pipeline has been demonstrated on well-studied benchmark Navier-Stokes simulations over a parameter range, each featuring a single bifurcation type, for which POD-based reductions have proven adequate. A natural next step is to showcase the pipeline’s efficiency in constructing global normal-form ROMs capable of capturing the full dynamics, including turbulent regimes, across broader parameter domains. The concept for the construction of “global” surrogate normal-forms via machine learning, but for low-dimensional systems of ODEs, can be traced back in the 2000s [101] using explainable machine learning. To deal with the high-dimensionality of the spatio-temporal Navier-Stokes equations, this approach can be extended by incorporating the invariant manifold theory [11, 12]. Such an approach has been recently proposed in [49] for dissipative 1D PDEs (the Chafee-Infante reaction-diffusion PDE and the Kuramoto-Sivashinsky PDE) using autoencoders, but also POD and Diffusion Maps, to parametrize an approximate inertial manifold (AIM) resulting in a “global” surrogate normal-form of ODEs. Its potential for application to the Navier-Stokes equations is strongly indicated by our recently results in [42], where we have shown that while the POD fails to identify the actual intrinsic dimension of the underlying latent space (i.e., the correct dimension of the normal-form model), Diffusion Maps succeed in identifying the appropriate dimension of the global dynamics.
Acknowledgements. The authors acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support. The PNRR MUR, projects PE0000013-Future Artificial Intelligence Research-FAIR & CN0000013 CN HPC - National Centre for HPC, Big Data and Quantum Computing, Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM) partially supported C.S.
Declaration of interests. The authors report no conflict of interest.